OpenStudy (anonymous):

if m∠HTE = 3x-10 and m∠HBE = x+40, what is m∠HTB?

OpenStudy (anonymous):



OpenStudy (anonymous):

According to your figure drawn. HTBE is a parallelogram. and ∠HTE and ∠HBE are its opposite angles. opposite angles of a parallelogram are congruent Hence m∠HTE = m∠HBE i.e. 3x-10 = x+40 i.e. 3x-x = 40+10 i.e. 2x = 50 i.e. x= 25  @Violetuserpurple